OpenStudy (anonymous):

The volume of a cylinder is 81 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@eliassaab @ganeshie8

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@ganeshie8 @UnkleRhaukus

OpenStudy (anonymous):

sorrry i don't know but ur queastion is terriffic

OpenStudy (anonymous):

it ok

OpenStudy (anonymous):

can you get someone to help me please

OpenStudy (gudden):



OpenStudy (mayankdevnani):

we know that :- \[\large \bf Vol.~of~ cone=\frac{1}{3}Vol.~of~ cylinder\] so, \[\large \bf Vol.~of~cylinder=3 \times Vol.~of~cone\]

OpenStudy (mayankdevnani):

now,can you solve it ? @prowrestler

OpenStudy (anonymous):

let me try

OpenStudy (mayankdevnani):

sure

OpenStudy (mayankdevnani):

just apply this formula :- \[\huge \bf Vol.~of~cone=\frac{1}{3} \times Vol.~of~cylinder\]

OpenStudy (mayankdevnani):

and plug all the values in that formula

OpenStudy (gudden):

@mayankdevnani .. i guess i said the same thing.. just a little over..ur comment..

OpenStudy (mayankdevnani):

yeah

OpenStudy (anonymous):

wait im doing it right now

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

so whats the volume of the cone and the cylinder

OpenStudy (gudden):

@prowrestler .. u just read my comments... n it would be clear to u (i guess)

OpenStudy (anonymous):

i gotcha

OpenStudy (anonymous):

im confused so whats the radius and the height i need to know them to plug it into the formuas

OpenStudy (gudden):

first of all see the diagram that i had drawn..

OpenStudy (anonymous):

and fan me guys

OpenStudy (anonymous):

yes i see it

OpenStudy (gudden):

so can u see that clearly the height of the cylinder and cone is equal..and so is the radius (its because cos, the question says that the cone "exactly" fits into the cylinder)

OpenStudy (anonymous):

sp what is it

OpenStudy (anonymous):

*so

OpenStudy (gudden):

since Volume of cylinder = pi r^2 h Volume of cone = 1/3 pi r^2 h and radius and height are equal for both so u can find out the value of volume of cone using \[volume of cone = \frac{ 1 }{ 3 } \pi r ^{2} h\]

OpenStudy (gudden):

and u know that here \[\pi r ^{2} h\] = 81

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thats the radius right

OpenStudy (gudden):

so is it clear now ??

OpenStudy (gudden):

n BTW which grade r u in?

OpenStudy (anonymous):

im 11th

OpenStudy (gudden):

ok..

OpenStudy (anonymous):

so whats the height for them both

OpenStudy (gudden):

thats same i guess it could be clear from the diagram i made.. and when it comes to radius. since u can see that they share.the common circular end.. so since the circumference is same therefore as a result the radius is also same...

OpenStudy (anonymous):

so thats the answer @gudden

OpenStudy (gudden):

27 cubic feet would come out to be ur final answer..

OpenStudy (anonymous):

Thnks man!!

OpenStudy (gudden):

i m a gal...xd,..

OpenStudy (anonymous):

oh sorry:)

OpenStudy (gudden):

its alright...:D